find the equation of parabola , when the vertex is(2,1) and directrix is x=y-1

Grade:12th pass

1 Answers

Arun
6 years ago
Dear student
ย 

Equation of directrix=ย x - y+1=0--------(i) ย 

ย 

Let the coordinate of focus = (a,b)

ย 

ย the axis of parabola is perpendicular to directrix.

ย 

So the equation of axis of parabola may be taken as x+y+k=0

ย 

it passes through ย (2,1)

โ‡’ 2+1+k=0

โ‡’ k=-3

ย 

โ‡’ the equation of axis of parabola is x+y-3=0---------------(ii)

ย 

Now, for the point of intersection of directrix and axis of parabola

On solving (i) and (ii), we get,

ย 

y = 2

put the value of y in (i), we get,ย 

x = 1

ย 

Thus, the point of intersection of directrix and axis of parabolaย 

(X, Y) = (1, 2)

Coordinates ofย 

z = (1,2)

ย 

Vertex A (2,1) is the mid-point of focus andย 

z = (1,2)

2 = (a+1)/2

a = 3

ย 

1 = (b+2)/2

b = 0

ย 

Coordinate of focus = (3,0)

ย 

ย 

Now you can simply get the equation of parabola by

PS = PM